OpenStudy (bossimbacon):

How many solutions does the equation have? 4(2x+3)=2(3x-4) A. one solution B. infinite solutions C. no solution

OpenStudy (bossimbacon):

@shahd-owies

OpenStudy (immanuelv):

well you can cross out C

OpenStudy (immanuelv):

and it would have infinite solution only if both solutions are the same. In this case its not! xD

OpenStudy (bossimbacon):

so its A?

OpenStudy (immanuelv):

yup xD

OpenStudy (bossimbacon):

thanks <3

OpenStudy (immanuelv):

you're welcome :)

OpenStudy (anonymous):

4(2x+3)=2(3x-4) 8x+12=6x-8 8x-6x=-8-12 2x=-20 x=-20/2=-10 so x=-10 so the answer is A